Solve the following equations:
(q/4) + 7 = 5........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing 7 from LHS to RHS it becomes -7
Then,
= q/4 = 5–7
= q/4 = - 2
Now,
Multiply both side by 4,
= (q/4) × 4= -2 × 4
= a = -8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
Q/4 + 7 = 5
Q/4 = 5-7
Q/4 = -2
Q = -8
 
Hope it helps!
Thank you ✌️